# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v3

@@ Abstract
Abstract: The processes of gene expression are inherently stochastic, even for essential genes required for growth. How does the cell maximize fitness in light of noise? To answer this question, we build a mathematical model to explore the trade-off between metabolic load and growth robustness. The model predicts novel principles of central dogma regulation: Optimal protein expression levels for many genes are in vast overabundance. Essential genes are transcribed above a lower limit of one message per cell cycle. Gene expression is achieved by load balancing between transcription and translation. We present evidence that each of these novel regulatory principles is observed. These results reveal that robustness and metabolic load determine the global regulatory principles that govern central dogma processes, and these principles have broad implications for cellular function.
@@ Introduction
Figure 1: The RLTO Model. Panel A: Gene expression is stochastic. The central dogma describes a two-stage stochastic process in which genes are first transcribed and then translated. The transcription process transcribes an average of $\mu_{m}$ messages per cell cycle. The translation process translates an average of $\varepsilon$ proteins per message. Panel B & C: A schematic cell lineage tree is shown during exponential growth. For a specific protein $i$, the cell fill represents the protein number $N_{p}$ relative to its threshold number $n_{p}$ required for cell growth. Panel B: Reducing the mean expression level reduces doubling time; however, stochasticity in expression results in below-threshold cells (red fill) which grow slowly. Panel C: Increasing protein expression increases the doubling time; however, all cells are above threshold (blue fill). Panel D: The fitness landscape as a function of protein number. Growth arrests for protein number $N_{p}$ smaller than the threshold level $n_{p}$ (red) due to the failure of essential processes. High expression levels are penalized due to the metabolic cost of protein expression. This trade-off leads to a highly asymmetric fitness landscape: The relative metabolic cost of overabundance is small relative to the cost of growth arrest due to the large size of the total metabolic load $N_{0}$. Panel E: The gene expression process is stochastic. There is significant cell-to-cell variation in protein abundance ($N_{p}$) around the mean level ($\mu_{p}$). Even for mean expression levels significantly above the threshold level $n_{p}$, some cells fall below threshold (red). The distribution in protein number is modeled using a gamma distribution [9], parameterized by message number $\mu_{m}$ and translation efficiency $\varepsilon$. Panel F: The robustness load trade-off determine the optimal expression level. The population growth rate depends on the distribution of the protein number. The asymmetry of the fitness landscape drives the optimal expression level far above the threshold level due to the high fitness cost of low protein abundance.
@@ Results
Defining the RLTO Model. To study the consequences of growth robustness on the central dogma quantitatively, we propose and analyze a minimal model: the Robustness-Load Trade-Off (RLTO) Model. The model includes three critical components: (i) Protein levels are stochastic and the single-cell growth rate depends upon them, (ii) gene transcription and translation generate a metabolic load, and (iii) cell growth is dependent on a large number of essential genes. These model characteristics result in a highly-asymmetric fitness landscape. The optimization of expression on this asymmetric landscape predicts new phenomenology absent from previous models (e.g. [10]).
@@
The protein number $N_{p}$ expressed from gene $i$ is the product of two sequential stochastic processes: transcription and translation [11], leading to cell-to-cell variation in protein number, which we will refer to as noise. In our analysis, we will model gene expression using the canonical equilibrium Paulsson model [12]. (See Fig. 1A.) In this model, the numbers of proteins $N_{p}$ for gene $i$ are predicted to be gamma-distributed [13], in close agreement with observation [9]. The distribution is described by two gene-specific statistical parameters: message number ($\mu_{m}$), defined as the mean number of messages transcribed per cell cycle for gene $i$, and the translation efficiency ($\varepsilon$), the mean number of proteins translated from each message transcribed for gene $i$. The mean protein abundance is their product: $\mu_{p}=\mu_{m}\varepsilon$. These parameters can be expressed in terms of ratios of the rates of the underlying central dogma processes, as described in Supplementary Material Sec. A.1.
@@
How should the effect of essential protein expression on growth rate be modeled in the context of the RLTO model? Much recent work has focused on cellular resource allocation to functional sectors (e.g. [14]). In this approach, an optimization is performed by coherently modulating the abundance of all protein in a particular sector, leading to a trade-off between functional capacities of the cell. However, in the RLTO model, the optimization is fundamentally different: We consider the incoherent variation in the abundance of protein species $i$ due to noise. For these incoherent changes, we generically expect proteins to exhibit rate-limited kinetics: Increases in the protein number $N_{p}$ above a threshold level $n_{p}$ has minimal effect on the rate since other chemical species (proteins, metabolites, etc.) are rate limiting [15]. However, if the protein number $N_{p}$ falls below the threshold $n_{p}$, then protein species $i$ becomes rate limiting and leads to a significant slowdown of the growth rate. In the RLTO model, we coarse-grain the details of this growth slowdown as growth arrest. (See Fig. 1BCD.) There is already some precedent for the use of this type of threshold (e.g. [16]), but we will demonstrate that the detailed form of the fitness landscape is not important. (See Supplementary Material Sec. B.) Although sufficiently detailed knowledge of the relevant molecular and cellular biology could be used to predict the protein thresholds $n_{p}$, we will treat these as gene-specific unknown parameters.
@@
where the first term represents fitness loss due to metabolic load of transcription and translation while the second term represents loss due to the arrest of essential processes. In summary, the model depends only on a single global parameter: the relative metabolic load $\Lambda$ and three gene-specific parameters: the threshold number $n_{p}$, the message number $\mu_{m}$ and the relative translation efficiency $\varepsilon/N_{0}$. We propose that the cell is regulated to maximize the growth rate with respect to transcription (message number) and translation (translation efficiency). The fitness landscape predicted by the RLTO model for representative parameters is shown in Fig. 2A.
Figure 2: The RLTO model predicts overabundance is optimal. Panel A: Fitness landscape determines optimal message number and translation efficiency. The fitness loss ($s\equiv\ln k_{\rm max}/k$) is shown as a function of message number ($\mu_{m}$) and translation efficiency ($\varepsilon$). The red dotted curve represents programs where the mean protein number is equal to the threshold ($\mu_{p}=n_{p}$) and the red dot represents the optimal regulatory program ($\hat{\mu}_{m}$, $\hat{\varepsilon}$). Panel B: Gene-expression noise. Due to the stochasticity of the central dogma processes at equilibrium, the protein number $N_{p}$ is gamma-distributed [12]. For high-expression genes, expression has low noise and the protein number is tightly distributed about its mean; however, for low-expression genes, expression is noisy and the distribution is extremely wide. Panel C: Overabundance is optimal for all genes. For high-expression genes, low overabundance is optimal ($\mu_{p}\approx n_{p}$); however, for low-expression genes, vast overabundance is optimal ($\mu_{p}\gg n_{p}$). From a quantitative perspective, overabundance depends on the relative load $\Lambda$; however, the qualitative dependence is invariant to over an orders-of-magnitude range of values.
@@
Understanding the rationale for overabundance. To explore both the robustness of the protein overabundance prediction and to understand its mathematical rationale, we explored a collection of more complex models numerically. (Supplementary Material Sec. B.) The key mathematical feature that drives overabundance is not the assumption of growth arrest, but rather the strong asymmetry of the fitness landscape: the high cost of protein underabundance and the low cost of protein overabundance. (See Fig. 1EF.) In the RLTO model, this asymmetry is parameterized by the relative load ($\Lambda$), defined as the relative metabolic cost of transcribing an additional message. Since we estimate that $\Lambda<10^{-5}$, this cost is very low relative to the total metabolic cost of the cell, therefore we expect this asymmetry, and the prediction of the RLTO model, to be robust.
@@
Message number distribution is conserved. To what extent is this human data consistent with the RLTO model? For human cells, our test of the one-message rule is too simplistic in two respects: (i) We ignore the significant transcriptional differences associated with distinct cell types and (ii) the essential gene classification itself is defined by the ability of mutants or knockdowns to proliferate in cell culture; in marked contrast to the in vivo context where cell proliferation is tightly regulated [20]. Due to these subtleties, we decided to take a complementary approach: We considered the distribution of three different transcriptional statistics for each gene: transcription rate, cellular message number, defined as the average number of messages instantaneously, and message number ($\mu_{m}$), defined as the number of messages transcribed in a cell cycle. (See Supplementary Material Sec. C.) The RLTO model predicts a one-message threshold with respect to message number, but not the other two statistics. We therefore predict that the message number distributions in each organism (E. coli, yeast, and human) should align for low expression genes with respect to message number, but not for the other two transcriptional statistics. Consistent with the predictions of the RLTO model, there is a striking alignment of message number for essential genes between all three model organisms and growth conditions for message number. (See Fig. 3D.) This alignment is non-trivial: It is not observed with respect to other transcriptional statistics (Fig. 3C and Supplementary Material Fig. 3). The observed alignment is consistent with a conserved transcriptional regulatory program from E. coli to human, and we hypothesize that the rationale for this alignment is the abutment of the lowest essential-gene transcription levels in each organism against the one-message threshold.
@@
In a typical biological context, $\mu_{m}\ll\varepsilon$ and as a result, noise production is dominated by the transcription step of the gene expression process [12, 13]. (A table of central dogma parameters for each model organism appears in the Supplementary Material Tab. S2.) Quantitatively, the Paulsson model predicts that the noise should be inversely related to the message number [12, 13]:
@@ Discussion
What are the biological implications of noise? Many important proposals have been made, including bet-hedging strategies, the necessity of feedback in gene regulatory networks, etc. [7]. Our model suggests that overcoming cell-to-cell variation may fundamentally reshape the metabolic budget: Typically, proteins constitute 50-60% of the dry mass of the cell [5] and therefore overabundance could increase the overall protein budget by a significant factor. Why does the cell tolerate this significant increase in metabolic load above what would be predicted by a resource allocation analysis (e.g. [24])? This program dramatically reduces the consequence of stochastic expression of proteins on the rate of single-cell proliferation.
@@
The principles that govern central dogma regulation. We propose that robustness to noise fundamentally shapes the central dogma regulatory program for all genes and predicts a number of key regulatory principles. (See Fig. 6.) For high-expression genes, load balancing implies that gene expression consists of both high-amplification translation and transcription. The resulting expression level has low overabundance relative to the threshold required for function. In contrast, for essential low-expression genes, a three-fold strategy is implemented: (i) overabundance raises the mean protein levels far above the threshold required for function, (ii) load balancing, and (iii) the one-message rule ensure that message number is sufficiently large to lower the noise of inherently-noisy, low-expression genes. We anticipate that these regulatory principles, in particular protein overabundance, will have important implications, not only for our understanding of central dogma regulation specifically, but for understanding the rationale for protein expression level and function in many biological processes.
@@
for a population of cells subject to stochastic arrest with probability $1-P_{+}$ per cell cycle where $k_{0}$ is the growth rate of the non-arrested cells. For each gene $i$, the Paulsson model predicts the protein number CDF in terms of message number $\mu_{m}$ and translation efficiency $\varepsilon$ [12]. Assuming the below-threshold probability is small, the probability that the cell is below threshold for gene $i$ is:
@@ A.1 Methods: Detailed description of the noise model
The equilibrium Paulsson model for the central dogma describes multiple steps in the gene expression process [12, 13, 9]: Transcription generates mRNA messages. These messages are then translated to synthesize the protein gene products [11]. Both mRNA and protein are subject to degradation and dilution [30]. At the single cell level, each of these processes are stochastic. We will model these processes with the stochastic kinetic scheme [11]:
@@
Consistent with previous reports [12, 13], we find that the distribution of protein number per cell (at cell birth) was described by a gamma distribution: $N_{p}\sim\Gamma(a,\theta)$, where $N_{p}$ is the protein number at cell birth and $\Gamma$ is the gamma distribution which is parameterized by a scale parameter $\theta$ and a shape parameter $a$. (See Sec. A.1.4.) We refer to this distribution as the Paulsson model; The relation between the four kinetic parameters and these two statistical parameters has already been reported, and have clear biological interpretations [13]: The scale parameter:
@@ A.1.3 Gillespie simulation of stochastic kinetic scheme
The gamma distributions for each mean message number with scale and shape parameters determined by the corresponding translation efficiency and message number ($\theta=\varepsilon\ln 2$, $a=\frac{\mu_{m}}{\ln 2}$) as used for the Gillespie simulation were also plotted with the protein distributions. We observe an excellent match between these Gillespie simulations and the Paulsson statistical noise model (i.e. gamma function) as shown in Fig. S1.

### v4

@@ Abstract
Abstract: The processes of gene expression are inherently stochastic, even for essential genes required for growth. How does the cell maximize fitness in light of noise? To answer this question, we build a mathematical model to explore the trade-off between metabolic load and growth robustness. The model predicts novel principles of central dogma regulation: Optimal protein expression levels for many genes are in vast overabundance. Essential genes are transcribed above a lower limit of one message per cell cycle. Gene expression is achieved by load balancing between transcription and translation. We present evidence that each of these novel regulatory principles is observed. These results reveal that robustness and metabolic load determine the global regulatory principles that govern gene expression processes, and these principles have broad implications for cellular function.
Short title: Noise robustness shapes central dogma regulation
@@ Introduction
Figure 1: The RLTO Model. Panel A: Gene expression is stochastic. The central dogma describes a two-stage stochastic process where genes are first transcribed and then translated. The transcription process transcribes an average of $\mu_{m}$ messages per cell cycle. The translation process translates an average of $\varepsilon$ proteins per message. Panel B & C: A schematic cell lineage tree is shown during exponential growth. For a specific protein $i$, the cell fill represents the protein number $N_{p}$ relative to its threshold number $n_{p}$ required for cell growth. Panel B: Reducing the mean expression level reduces doubling time; however, stochasticity in expression results in below-threshold cells (red fill) which grow slowly. Panel C: Increasing protein expression increases the doubling time; however, all cells are above threshold (blue fill). Panel D: The fitness landscape as a function of protein number. Growth arrests for protein number $N_{p}$ smaller than the threshold level $n_{p}$ (red) due to the failure of essential processes. High expression levels are penalized due to the metabolic cost of protein expression. This trade-off leads to a highly asymmetric fitness landscape: The relative metabolic cost of overabundance is small relative to the cost of growth arrest due to the large size of the total metabolic load $N_{0}$. Panel E: The gene expression process is stochastic. There is significant cell-to-cell variation in protein abundance ($N_{p}$) around the mean level ($\mu_{p}$). Even for mean expression levels significantly above the threshold level $n_{p}$, some cells fall below threshold (red). The distribution in protein number is modeled using a gamma distribution [9], parameterized by message number $\mu_{m}$ and translation efficiency $\varepsilon$. Panel F: The robustness load trade-off determines the optimal expression level. The population growth rate depends on the distribution of the protein number. The asymmetry of the fitness landscape drives the optimal expression level far above the threshold level due to the high fitness cost of low protein abundance.
@@ Results
Defining the RLTO Model. To study the consequences of growth robustness on gene expression quantitatively, we propose and analyze a minimal model: the Robustness-Load Trade-Off (RLTO) Model. The model includes three critical components: (i) Protein levels are stochastic and the single-cell growth rate depends upon them, (ii) gene transcription and translation generate a metabolic load, and (iii) cell growth is dependent on a large number of essential genes. These model characteristics result in a highly-asymmetric fitness landscape. The optimization of expression on this asymmetric landscape predicts new phenomenology absent from previous models (e.g. [10]).
@@
The protein number $N_{p}$ expressed from gene $i$ is the product of two sequential stochastic processes: transcription and translation [11], leading to cell-to-cell variation in protein number, which we will refer to as noise. In our analysis, we will model gene expression using the canonical steady-state noise model [12]. (See Fig. 1A.) In this model, the numbers of proteins $N_{p}$ for gene $i$ are predicted to be gamma-distributed [13], in close agreement with observation [9]. The distribution is described by two gene-specific statistical parameters: message number ($\mu_{m}$), defined as the mean number of messages transcribed per cell cycle for gene $i$, and the translation efficiency ($\varepsilon$), the mean number of proteins translated from each message transcribed for gene $i$. The mean protein abundance is their product: $\mu_{p}=\mu_{m}\varepsilon$. These parameters can be expressed in terms of ratios of the rates of the underlying gene expression processes, as described in Supplementary Material Sec. A.1.
@@
How should the effect of essential protein expression on growth rate be modeled in the context of the RLTO model? Much recent work has focused on cellular resource allocation to functional sectors (e.g. [14]). In this approach, an optimization is performed by the coordinated modulation of the abundance of all proteins in a particular sector, leading to a trade-off between functional capacities of the cell. However, in the RLTO model, the optimization is fundamentally different: We consider the uncoordinated modulation of the abundance of protein species $i$ due to noise. For these incoherent changes, we generically expect proteins to exhibit rate-limited kinetics: Increases in the protein number $N_{p}$ above a threshold level $n_{p}$ has minimal effect on the rate since other chemical species (proteins, metabolites, etc.) are rate limiting [15]. However, if the protein number $N_{p}$ falls below the threshold $n_{p}$, then protein species $i$ becomes rate limiting and leads to a significant slowdown of the growth rate. In the RLTO model, we coarse-grain the details of this growth slowdown as growth arrest. (See Fig. 1.) There is already some precedent for the use of this type of threshold (e.g. [16]), but we will demonstrate that the detailed form of the fitness landscape is not important. (See Supplementary Material Sec. B.) Although sufficiently detailed knowledge of the relevant molecular and cellular biology could be used to predict the protein thresholds $n_{p}$, we will treat these as gene-specific unknown parameters.
@@
where the first term represents fitness loss due to metabolic load of transcription and translation while the second term represents loss due to the arrest of essential processes. $\gamma$ is the regularized incomplete gamma function and the central distribution function (CDF) of the gamma distribution. (See Supplementary Material Sec. A.1 and  A.2.) In summary, the model depends only on a single global parameter: the relative metabolic load $\Lambda$ and three gene-specific parameters: the threshold number $n_{p}$, the message number $\mu_{m}$ and the relative translation efficiency $\varepsilon/N_{0}$. We propose that the cell is regulated to maximize the growth rate with respect to transcription (message number) and translation (translation efficiency). The fitness landscape predicted by the RLTO model for representative parameters is shown in Fig. 2A.
Figure 2: The RLTO model predicts overabundance is optimal. Panel A: Fitness landscape determines optimal message number and translation efficiency. The fitness loss ($s\equiv\ln k_{\rm max}/k$) is shown as a function of message number ($\mu_{m}$) and translation efficiency ($\varepsilon$). The red dotted curve represents programs where the mean protein number is equal to the threshold ($\mu_{p}=n_{p}$) and the red dot represents the optimal regulatory program ($\hat{\mu}_{m}$, $\hat{\varepsilon}$). Panel B: Gene-expression noise. Due to the stochasticity of gene expression processes at equilibrium, the protein number $N_{p}$ is gamma-distributed [9]. For high-expression genes, expression has low noise and the protein number is tightly distributed around its mean; however, for low-expression genes, expression is noisy and the distribution is extremely wide. Panel C: Overabundance is optimal for all genes. For high-expression genes, low overabundance is optimal ($\mu_{p}\approx n_{p}$); however, for low-expression genes, vast overabundance is optimal ($\mu_{p}\gg n_{p}$). From a quantitative perspective, overabundance depends on the relative load $\Lambda$; however, the qualitative dependence is invariant to over an orders-of-magnitude range of values.
@@
Understanding the rationale for overabundance. To explore both the robustness of the protein overabundance prediction and to understand its mathematical rationale, we explored a collection of more complex models numerically. (Supplementary Material Sec. B.) The key mathematical feature that drives overabundance is not the assumption of growth arrest, but rather the strong asymmetry of the fitness landscape: the high cost of protein underabundance and the low cost of protein overabundance. (See Fig. 1EF.) The population growth rate (Panel 1F) can be understood qualitatively as the convolution of the single-cell growth rate (Panel 1D) with the probability density function (PDF) of the protein abundance (Panel 1E). In the RLTO model, this asymmetry is parameterized by the relative load ($\Lambda$), defined as the relative metabolic cost of transcribing an additional message. Since we estimate that $\Lambda<10^{-5}$, this cost is very low relative to the total metabolic cost of the cell, therefore we expect this asymmetry, and the prediction of the RLTO model, to be robust.
@@
Message number distribution is conserved. To what extent is this human data consistent with the RLTO model? For human cells, our test of the one-message rule is too simplistic in two respects: (i) We ignore the significant transcriptional differences associated with distinct cell types and (ii) the essential gene classification itself is defined by the ability of mutants or knockdowns to proliferate in cell culture; in marked contrast to the in vivo context where cell proliferation is tightly regulated [20]. Due to these subtleties, we decided to take a complementary approach: We considered the distribution of three different transcriptional statistics for each gene: transcription rate, cellular message number, defined as the average number of messages instantaneously, and message number ($\mu_{m}$), defined as the number of messages transcribed in a cell cycle. (See Supplementary Material Sec. C.) The RLTO model predicts a one-message threshold with respect to message number, but not the other two statistics. We therefore predict that the message number distributions in each organism (E. coli, yeast, and human) should align for low expression genes with respect to message number, but not for the other two transcriptional statistics. Consistent with the predictions of the RLTO model, there is a striking alignment of message number for essential genes between all three model organisms and growth conditions for message number. (See Fig. 3D.) This alignment is non-trivial: It is not observed with respect to other transcriptional statistics (Fig. 3C and Supplementary Material Fig. 3).
What is the significance of the similarity in the distributions of message number between organisms? Another strategy for satisfying the one message rule would be for transcription to be increased. For instance, mammalian cells have about 1000 times the number of proteins relative to bacterial cells. (See Supplemental Tab. S2.) One might therefore naively predict that the message number should be increased 1000-fold as well. This is not observed. In fact, the message number distributions of all the model organisms analyzed abut the one message threshold. The proximity to the threshold suggests that organisms do as little transcription as possible while satisfying the one message rule. This appears to be a conserved transcriptional regulatory strategy from E. coli to human.
@@
In a typical biological context, $\mu_{m}\ll\varepsilon$ and as a result, noise production is dominated by the transcription step of the gene expression process [12, 13]. (A table of central dogma parameters for each model organism appears in the Supplementary Material Tab. S2.) Quantitatively, the canonical steady-state noise model predicts that the noise should be inversely related to the message number [12, 13]:
@@ Discussion
What are the biological implications of noise? Many important proposals have been made, including bet-hedging strategies, the necessity of feedback in gene regulatory networks, etc. [7]. Our model suggests that overcoming cell-to-cell variation may fundamentally reshape the metabolic budget: Typically, proteins constitute 50-60% of the dry mass of the cell [5] and therefore overabundance could increase the overall protein budget by a significant factor. Why does the cell tolerate this significant increase in metabolic load above what would be predicted by a resource allocation analysis (e.g. [24])? This strategy dramatically reduces the consequence of stochastic expression of proteins on the rate of single-cell proliferation.
@@
The principles that govern central dogma regulation. We propose that robustness to noise fundamentally shapes the central dogma regulatory program for all genes and predicts a number of key regulatory principles. (See Fig. 6.) For high-expression genes, load balancing implies that gene expression consists of both high-amplification translation and transcription. The resulting expression level has low overabundance relative to the threshold required for function. In contrast, for essential low-expression genes, a three-fold strategy is implemented: (i) overabundance raises the mean protein levels far above the threshold required for function, (ii) load balancing, and (iii) the one-message rule ensures that message number is sufficiently large to lower the noise of inherently-noisy, low-expression genes. We anticipate that these regulatory principles, in particular protein overabundance, will have important implications, not only for our understanding of central dogma regulation specifically, but for understanding the rationale for protein expression level and function in many biological processes.
@@
for a population of cells subject to stochastic arrest with probability $1-P_{+}$ per cell cycle where $k_{0}$ is the growth rate of the non-arrested cells. For each gene $i$, the canonical steady-state noise model predicts the protein number CDF in terms of message number $\mu_{m}$ and translation efficiency $\varepsilon$ [12]. Assuming the below-threshold probability is small, the probability that the cell is below threshold for gene $i$ is:
@@ A.1 Methods: Detailed description of the noise model
The canonical steady-state noise model for the central dogma describes multiple steps in the gene expression process [12, 13, 9]: Transcription generates mRNA messages. These messages are then translated to synthesize the protein gene products [11]. Both mRNA and protein are subject to degradation and dilution [30]. At the single cell level, each of these processes are stochastic. We will model these processes with the stochastic kinetic scheme [11]:
@@
Consistent with previous reports [12, 13], we find that the distribution of protein number per cell (at cell birth) was described by a gamma distribution: $N_{p}\sim\Gamma(a,\theta)$, where $N_{p}$ is the protein number at cell birth and $\Gamma$ is the gamma distribution which is parameterized by a scale parameter $\theta$ and a shape parameter $a$. (See Sec. A.1.4.) We refer to this distribution as the canonical steady-state noise model; The relation between the four kinetic parameters and these two statistical parameters has already been reported, and have clear biological interpretations [13]: The scale parameter:
@@ A.1.3 Gillespie simulation of stochastic kinetic scheme
The gamma distributions for each mean message number with scale and shape parameters determined by the corresponding translation efficiency and message number ($\theta=\varepsilon\ln 2$, $a=\frac{\mu_{m}}{\ln 2}$) as used for the Gillespie simulation were also plotted with the protein distributions. We observe an excellent match between these Gillespie simulations and the canonical steady-state statistical noise model (i.e. gamma function) as shown in Fig. S1.